# Privately Evaluating Untrusted Black-Box Functions

### v1

@@ 1 Introduction
In addition to addressing the black-box privacy problem, our private mechanisms provide feasibility results for differentially private release of general classes of functions. Most work on differential privacy considers the easier, “white box” setting, when a complete description of the input function $f$ is available. Our results have applications in this setting. We discuss this perspective on our results in Section 1.1.4.
@@ 1.1.1 Automated Sensitivity Detection
The wrapper we design, Sens-o-Matic, is private for all functions $f:\mathcal{U}^{*}\to\mathcal{Y}$ with finite range $\mathcal{Y}$ and extends the accuracy guarantees of (both versions of) ShI from monotone functions to all functions. It works by running ShI (or our modification thereof) on a carefully selected “monotonization” of function $f$—see Section 1.2 for more detail. To complete the picture for automated sensitivity detection setting, we provide a lower bound (in Corollary 5.2) on the locality $\lambda$ of any privacy wrapper that achieves accuracy equal to down sensitivity at depth $\lambda$. Our lower bound matches the first term (corresponding to the pure DP) in the guarantee for Sens-o-Matic and nearly matches the second term (corresponding to approximate DP). We also provide a query complexity lower bound, Theorem 6.1 (and Remark 6.2), showing that the query complexity we achieve cannot be significantly improved, even if the locality requirement is relaxed. Our guarantees for Sens-o-Matic are compared to guarantees for both variants of ShI and the locality lower bound in Table 1.
@@ 1.1.2 Privacy Wrappers with Claimed Sensitivity Bound
Sensitivity and the related notion of Lipschitz functions play a fundamental role in private data analysis. Intuitively, sensitivity measures how small modifications of the dataset affect the value of the function. Given a constant $c>0$ and a domain $D\subseteq\mathcal{U}^{*}$, a function $f:\mathcal{U}^{*}\to\mathbb{R}$ is $c$-Lipschitz over $D$ if $|f(x)-f(y)|\leq c$ for all neighbors $x,y\in D$. The smallest constant $c$ for which function $f$ is $c$-Lipschitz on $\mathcal{U}^{*}$ is called the Lipschitz constant or the (global) sensitivity of $f$. A basic result is that the Laplace mechanism, which releases $f(x)$ with Laplace noise scaled to $\frac{c}{\varepsilon}$, is ($\varepsilon,0)$-differentially private [DMNS06]. Most work on differential privacy considers the “white box” case, where a complete description of $f$ is available, and one analyzes sensitivity analytically.
The problem of privately evaluating black-box functions was first considered by Jha and Raskhodnikova [JR13]. In their setting, in addition to the black-box function $f$, the analyst provides [^3] Equivalently, the analyst provides a rescaled function $f/c$ instead of $f$, and the curator presumes a sensitivity bound of 1. a sensitivity parameter $c$. The data curator must guarantee differential privacy whether or not the provided bound is correct, but the mechanism’s accuracy depends on the bound being correct and as tight as possible. Jha and Raskhodnikova [JR13] and follow-up work [AJMR15, LLRV25] handle the case where the data universe $\mathcal{U}$ is finite, and they investigate the function’s behavior in a ball around the input that includes both insertions and deletions. This means the data universe has to be small to achieve reasonable query complexity, and also that the function must be robust to both insertions and deletions of data points.
@@
Lower Bounds [GRS09] and $(\varepsilon,\delta)$-DP $c$-Lipschitz $\dfrac{c}{\varepsilon}$ $\Omega{\Big{(}{\dfrac{1}{\varepsilon}\log\dfrac{1}{\delta}}\Big{)}}$
Thms. 5.1,6.1 [GRS09] Thms. 5.1,6.1
@@
We provide a novel privacy wrapper, Subset-Extension, for this setting with accuracy and locality guarantees that are each optimal up to constant factors. Table 2 summarizes our results and compares them to the guarantees of existing privacy mechanisms that query $f$ only on the subsets of dataset $x$, which we summarize briefly here. In particular, we provide a self-contained accuracy analysis of TAHOE in Proposition B.1. It adds Laplace noise that is larger by a factor of $\log(1/\delta)/\varepsilon$ than the scale $c/\varepsilon$ that can be used in the white-box setting when the function $f$ is promised to be $c$-Lipschitz on its entire domain. (This latter scale is known to be optimal, even for the special case where $f$ is a simple counting query [GRS09].)
@@
Our Subset-Extension privacy wrapper, whose performance is stated in Theorem 4.1, achieves the best of both worlds: the accuracy of Cummings-Durfee and the locality of TAHOE. We describe its construction, which departs significantly from the existing approaches, in Section 1.2. Both the error scale and locality of our privacy wrappers are essentially optimal. The optimality of error follows the work of Ghosh et al. [GRS09], mentioned above. To prove optimality of locality, we give two lower bounds: one which bounds locality directly (Theorem 5.1), and the other which bounds query complexity (Theorem 6.1), and thus locality by implication.
@@ 1.1.3 Privacy Wrappers with Claimed Sensitivity Bound for Bounded-Range Functions
In Appendices C and D, we show that one can achieved improved guarantees in the claimed sensitivity bound setting for functions with small intended range. We present our results for this setting in Table 3.
@@
Double Monotonization Thm. D.1 on ${\mathcal{N}}^{\downarrow}_{\lambda}(x)$ $O{\Big{(}{\dfrac{c}{\varepsilon}}\Big{)}}$ $O{\left({\dfrac{1}{\varepsilon}\log\dfrac{r}{c\beta}}\right)}$
Lower Bounds [GRS09] and $c$-Lipschitz $\dfrac{c}{\varepsilon}$ $\tilde{\Omega}{\Big{(}{\dfrac{1}{\varepsilon}\log\dfrac{\varepsilon r}{c\beta}}\Big{)}}$
Thms. 5.1 and 6.1 [GRS09] Thms. 5.1 and 6.1
The Double Monotonization mechanism extends the guarantees of Subset-Extension to the setting of bounded range, replacing the $\log(1/\delta)$ factor in the locality with $\log r$. At a technical level, it builds on Sens-o-Matic, modifying it to take advantage of the (claimed) Lipschitz constant provided by the analyst. In contrast, the Small Diameter Subset Extension mechanism is based on a novel approach to local Lipschitz filters. This approach follows the spirit of [JR13], who initiated an approach based on sublinear-time algorithms that was later studied in [AJMR15, LLRV25]. As a corollary of our techniques, we construct a local Lipschitz filter (Corollary C.2) that improves on [LLRV25]. The resulting privacy wrapper achieves the best of both the algorithms of [LLRV25], and [CD20].
@@ 1.1.4 Applications of Privacy Wrappers to White-Box Setting
Private mean estimation: As a simple illustration of our results, we show how they recover known results [NRS07, DL09] on private mean estimation in one dimension. If the dataset $x\in\mathbb{R}^{n}$ is contained in an unknown interval of radius $\sigma$, then applying each of our wrappers to the average function leads to privately releasing $\mu$ with error roughly $\sigma/\varepsilon n$. See Corollary A.1.
Empirical risk minimization: In Section A.2, we improve upon the user-level (or person-level) private empirical risk minimization algorithm of [GKK+23a] for the setting of one dimensional parameter spaces. In particular, in Theorems A.4 and A.3, we show that in empirical risk minimization for a one-dimensional parameter space $\mathcal{Y}$, the dependence on privacy parameters $\varepsilon$ and $\delta$ can be reduced from $O{\left({\log(1/\delta)^{2}/\varepsilon^{5/2}}\right)}$ to the minimum of $O{\left({\log(1/\delta)/\varepsilon^{3/2}}\right)}$ and $(1/\varepsilon)\log(1/\delta)\exp(O\log^{*}|\mathcal{Y}|)$. The first term in the minimum follows immediately by substituting the Subset-Extension mechanism for the relevant subroutine in the private empirical risk minimization algorithm of [GKK+23a]. The second term leverages the Sens-o-Matic mechanism, as well as a straightforward extension of one of the key tools used in [GKK+23a] to bound the Lipschitz constant on ${\mathcal{N}}^{\downarrow}_{\lambda}(x)$ for $\lambda\geq|x|/2$.
Estimating graph parameters: Our results give simpler derivations of the rate at which one can estimate the parameter $p$ of an Erdos-Renyi graph model $G(n,p)$ subject to node privacy guarantees. There is a node-private algorithm [BCSZ18b] that, for all $p$, given a graph drawn from $G(n,p)$, generates an estimate $\hat{p}$ with additive error $\frac{1}{n}+O{\big{(}{\frac{\sqrt{\log n}}{\varepsilon n^{3/2}}}\big{)}}$ for $p$ bounded away from 0 and 1 (efficient algorithms achieving a similar bound were later given by [SU21, CDHS24]). This bound was surprising since the non-private estimator is a simple sum whose local sensitivity, which is $\Theta(1/n)$ on typical graphs from $G(n,p)$, is too large to obtain the optimal rate by simply applying the Laplace mechanism. Such a strategy would lead to error $1/(\varepsilon n)$, instead of $1/(\varepsilon n^{3/2})$.
@@ 1.3 Related Work
Privacy in the context of black-box functions was first explicitly considered by [JR13]. They connect the claimed sensitivity setting with the concept of local filter from the sublinear algorithms literature. This line of work [JR13, AJMR15, LLRV25] constructs privacy wrappers from local filters for Lipschitz functions. Their constructions query the function $f$ on inputs obtained by both insertions and deletions. Their query complexity depends on the universe size, and the analyst’s sensitivity bound must allow for the insertion of arbitrary outliers in the dataset.
@@
One can interpret the information-theoretic results of [SCV18] as a generic transformation that takes a function $f$ and produces a version that is robust in the neighborhood of an input $x$ whenever $x$ is resilient with respect to the original function $f$. However, we are not aware of a generic way to transform that into a differentially private mechanism using previous work, without still needing to explore the values of $f(z)$ for all sets $z$ that differ from $x$ in $\lambda$ insertions and deletions. In Appendix E, we make the [SCV18] transformation explicit and discuss its consequences.
@@
[DL09] and [Smi11] give generic transformations that create differentially private versions of statistical estimators. These results are incomparable to ours, since they rely on specific properties of the estimators, such as asymptotic normality and low bias.
@@ 2 Preliminaries
Definition 2.1 (Neighboring sets, differential privacy [DMNS06]).
@@
When $\delta=0,$ this guarantee is referred to as pure differential privacy, and the guarantee with $\delta>0$ is called approximate differential privacy.
@@
Fact 2.1 (Laplace mechanism [DMNS06]).
@@
Fix $\varepsilon>0$ and $\delta\in(0,1)$. Let ${\mathcal{A}}$ be any algorithm and suppose ${\mathcal{M}}$ is an $(\varepsilon,\delta)$-DP mechanism. Then the algorithm that, on input $x$, runs ${\mathcal{A}}$ on the output of ${\mathcal{M}}(x)$, is $(\varepsilon,\delta)$-DP.
@@ 3 Privacy Wrappers with Automated Sensitivity Detection
For every universe $\mathcal{U}$, privacy parameters $\varepsilon>0$ and $\delta\in[0,1)$, correctness parameter $\beta\in[0,1)$, and finite range $\mathcal{Y}\subset\mathbb{R}$ with size $k\stackrel{{\scriptstyle{\mbox{\tiny def}}}}{{=}}|\mathcal{Y}|$, there exist a $\lambda$-down-local privacy wrapper $\mathcal{W}$ such that for every function $f:\mathcal{U}^{*}\to\mathcal{Y}$ and dataset $x\in\mathcal{U}^{*}$, with probability at least $1-\beta$,
@@ 3.1 Shifted Inverse Mechanism: Promised Monotone Functions
In this section, we describe a privacy mechanism that works under a promise that the black-box function $f$ is monotone. It can be viewed as a privacy wrapper under the promise, but we do not call it a privacy wrapper to stress that, in contrast to our privacy wrappers, this mechanism is not private when the promise is broken. Our mechanism is a variant [^4] See Footnote 2. of the Shifted Inverse (ShI) mechanism of Fang, Dong, and Yi [FDY22]. The original ShI mechanism satisfies pure differential privacy for all monotone functions $f$. Our variant gets better dependence on the size of the range of function $f$ at the expense of providing only approximate differential privacy.
@@
Now we are ready to present our generalization of the ShI mechanism from [FDY22]. We begin by defining the inverse loss function: Given a function $f:\mathcal{U}^{*}\to\mathbb{R}$, we define $\ell^{f}:\mathcal{U}^{*}\times\mathbb{R}\to[0,\infty]$ by [^5] There is an annoying technicality: If $y<f(\emptyset)$, then $\ell^{f}(x,y)=\min\emptyset=+\infty$.
@@
The advantage of the inverse loss function is that it has low sensitivity, even when $f$ has high sensitivity. However, this only holds when $f$ is monotone.
@@
Now observe that $g(x,j)=1\iff\ell^{f}(x,y_{j})\geq\lambda+1$. That is, the exact value of $\ell^{f}(x,y)$ does not matter past the threshold $\lambda+1$. So we do not need to compute the exact value of $\ell^{f}(x,y)$ in this case. Therefore, we can compute $g(x,j)$ using only the values of $f(s)$ on $s\in{\mathcal{N}}^{\downarrow}_{\lambda}(x)$. In symbols, we have
$g(x,j)$ $=\max\Big{\{}0,1-\frac{1}{\lambda+1}\ell^{f}(x,y_{j})\Big{\}}$
$=\max\Big{\{}0,1-\frac{1}{\lambda+1}\min\{|x\setminus s|:s\subseteq x,f(s)\leq y\}\Big{\}}$
$=\max\Big{\{}0,1-\frac{1}{\lambda+1}\min\left(\{\lambda+1\}\cup\{|x\setminus s|:s\subseteq x,f(s)\leq y\}\right)\Big{\}}$
$=\max\Big{\{}0,1-\frac{1}{\lambda+1}\min\left(\{\lambda+1\}\cup\{|x\setminus s|:s\in{\mathcal{N}}^{\downarrow}_{\lambda}(x),f(s)\leq y\}\right)\Big{\}}.$
3.2 Sens-o-Matic: Removing the Monotonicity Promise

### v2

@@ 1 Introduction
In addition to addressing the black-box privacy problem, our private mechanisms provide feasibility results for differentially private release of general classes of functions. Most work on differential privacy considers the easier “white box” setting, when a complete description of the input function $f$ is available. Our results have applications in this setting. We discuss this perspective on our results in Section 1.1.4.
@@ 1.1.1 Automated Sensitivity Detection
The wrapper we design, Sens-o-Matic, is private for all functions $f:\mathcal{U}^{*}\to\mathcal{Y}$ with finite range $\mathcal{Y}$ and extends the accuracy guarantees of (both versions of) ShI from monotone functions to all functions. It works by running ShI (or our modification thereof) on a carefully selected “monotonization” of function $f$—see Section 1.2 for more detail.
To complete the picture for automated sensitivity detection setting, we provide a lower bound (in Corollary 5.2)  on the locality $\lambda$ of any privacy wrapper that achieves accuracy equal to down sensitivity at depth $\lambda$. Our lower bound matches the first term (corresponding to the pure DP) in the guarantee for Sens-o-Matic and nearly matches the second term (corresponding to approximate DP). We also provide a query complexity lower bound, Theorem 6.1 (and Remark 6.2), showing that the query complexity we achieve cannot be significantly improved, even if the locality requirement is relaxed. Our guarantees for Sens-o-Matic are compared to guarantees for both variants of ShI and the locality lower bound in Table 1.
@@ 1.1.2 Privacy Wrappers with Claimed Sensitivity Bound
Sensitivity and the related notion of Lipschitz functions play a fundamental role in private data analysis. Intuitively, sensitivity measures how small modifications of the dataset affect the value of the function. Given a constant $c>0$ and a domain $D\subseteq\mathcal{U}^{*}$, a function $f:\mathcal{U}^{*}\to\mathbb{R}$ is $c$-Lipschitz over $D$ if $|f(x)-f(y)|\leq c$ for all neighbors $x,y\in D$. The smallest constant $c$ for which function $f$ is $c$-Lipschitz on $\mathcal{U}^{*}$ is called the Lipschitz constant or the (global) sensitivity of $f$. A basic result is that the Laplace mechanism, which releases $f(x)$ with Laplace noise scaled to $\frac{c}{\varepsilon}$, is ($\varepsilon,0)$-differentially private [DMNS16]. Most work on differential privacy considers the “white box” case, where a complete description of $f$ is available, and one analyzes sensitivity analytically.
The problem of privately evaluating black-box functions was first considered by Jha and Raskhodnikova [JR13]. In their setting, in addition to the black-box function $f$, the analyst provides [^3] Equivalently, the analyst provides a rescaled function $f/c$ instead of $f$, and the curator presumes a sensitivity bound of 1. a sensitivity parameter $c$. The data curator must guarantee differential privacy whether or not the provided bound is correct, but the mechanism’s accuracy depends on the bound being correct and as tight as possible. Jha and Raskhodnikova [JR13] and follow-up work [AJMR14, LLRV25] handle the case where the data universe $\mathcal{U}$ is finite, and they investigate the function’s behavior in a ball around the input that includes both insertions and deletions. This means the data universe has to be small to achieve reasonable query complexity, and also that the function must be robust to both insertions and deletions of data points.
@@
Lower Bounds [GRS12] and $(\varepsilon,\delta)$-DP $c$-Lipschitz $\dfrac{c}{\varepsilon}$ $\Omega{\Big{(}{\dfrac{1}{\varepsilon}\log\dfrac{1}{\delta}}\Big{)}}$
Thms. 5.1,6.1 [GRS12] Thms. 5.1,6.1
@@
We provide a novel privacy wrapper, Subset-Extension, for this setting with accuracy and locality guarantees that are each optimal up to constant factors. Table 2 summarizes our results and compares them to the guarantees of existing privacy mechanisms that query $f$ only on the subsets of dataset $x$, which we summarize briefly here. In particular, we provide a self-contained accuracy analysis of TAHOE in Proposition B.1. It adds Laplace noise that is larger by a factor of $\log(1/\delta)/\varepsilon$ than the scale $c/\varepsilon$ that can be used in the white-box setting when the function $f$ is promised to be $c$-Lipschitz on its entire domain. (This latter scale is known to be optimal, even for the special case where $f$ is a simple counting query [GRS12].)
@@
Our Subset-Extension privacy wrapper, whose performance is stated in Theorem 4.1, achieves the best of both worlds: the accuracy of Cummings-Durfee and the locality of TAHOE. We describe its construction, which departs significantly from the existing approaches, in Section 1.2.
Both the error scale and locality of our privacy wrappers are essentially optimal. The optimality of error follows the work of Ghosh et al. [GRS12], mentioned above. To prove optimality of locality, we give two lower bounds: one which bounds locality directly (Theorem 5.1), and the other which bounds query complexity (Theorem 6.1), and thus locality by implication.
@@ 1.1.3 Privacy Wrappers with Claimed Sensitivity Bound for Bounded-Range Functions
In Appendices C and D, we show that one can achieve improved guarantees in the claimed sensitivity bound setting for functions with small intended range. We present our results for this setting in Table 3.
@@
Double Monotonization Thm. D.1 on ${\mathcal{N}}^{\downarrow}_{\lambda}(x)$ $O{\left({\dfrac{1}{\varepsilon}\log\dfrac{r}{c\beta}}\right)}$
Lower Bounds [GRS12] and $c$-Lipschitz $\dfrac{c}{\varepsilon}$ $\tilde{\Omega}{\Big{(}{\dfrac{1}{\varepsilon}\log\dfrac{\varepsilon r}{c\beta}}\Big{)}}$
Thms. 5.1 and 6.1 [GRS12] Thms. 5.1 and 6.1
The Double Monotonization mechanism extends the guarantees of Subset-Extension to the setting of bounded range, replacing the $\log(1/\delta)$ factor in the locality with $\log r$. At a technical level, it builds on Sens-o-Matic, modifying it to take advantage of the (claimed) Lipschitz constant provided by the analyst.
In contrast, the Small Diameter Subset Extension mechanism is based on a novel approach to local Lipschitz filters. This approach follows the spirit of [JR13], who initiated an approach based on sublinear-time algorithms that was later studied in [AJMR14, LLRV25]. As a corollary of our techniques, we construct a local Lipschitz filter (Corollary C.2) that improves on [LLRV25]. The resulting privacy wrapper achieves the best of both the algorithms of [LLRV25], and [CD20].
@@ 1.1.4 Applications of Privacy Wrappers to White-Box Setting
Private mean estimation: As a simple illustration of our results, we show how they recover known bounds [NRS07, DL09] on private mean estimation in one dimension. If the dataset $x\in\mathbb{R}^{n}$ is contained in an unknown interval of radius $\sigma$, then applying each of our wrappers to the average function leads to privately releasing $\mu$ with error roughly $\frac{\sigma}{\varepsilon n}$. See Corollary A.1 for details.
Empirical risk minimization: In Section A.2, we improve upon the user-level (also known as person-level) private empirical risk minimization algorithm of [GKK+23a] for the setting of one dimensional parameter spaces. In particular, in Theorems A.4 and A.3, we show that in empirical risk minimization for a one-dimensional parameter space $\mathcal{Y}$, the dependence on privacy parameters $\varepsilon$ and $\delta$ can be reduced from $O\big{(}\frac{1}{\varepsilon^{5/2}}\log^{2}\frac{1}{\delta}\big{)}$ to the minimum of  $O\big{(}\frac{1}{\varepsilon^{3/2}}\log\frac{1}{\delta}\big{)}$ and $\frac{1}{\varepsilon}\big{(}\log\frac{1}{\delta}\big{)}\exp(O\log^{*}|\mathcal{Y}|)$. The first term in the minimum follows immediately by substituting the Subset-Extension mechanism for the relevant subroutine in the private empirical risk minimization algorithm of [GKK+23a]. The second term leverages the Sens-o-Matic mechanism, as well as a straightforward extension of one of the key tools used in [GKK+23a] to bound the Lipschitz constant on ${\mathcal{N}}^{\downarrow}_{\lambda}(x)$ for $\lambda\geq|x|/2$.
Estimating graph parameters: Our results give simpler derivations of the rate at which one can estimate the parameter $p$ of an Erdős–Rényi graph model $G(n,p)$ subject to node privacy guarantees. There is a node-private algorithm [BCSZ18b] that, for all $p$, given a graph drawn from $G(n,p)$, generates an estimate $\hat{p}$ with additive error $\frac{1}{n}+O{\big{(}{\frac{\sqrt{\log n}}{\varepsilon n^{3/2}}}\big{)}}$ for $p$ bounded away from 0 and 1 (efficient algorithms achieving a similar bound were later given by [SU21, CDHS24]). This bound was surprising since the non-private estimator is a simple sum whose local sensitivity, which is $\Theta(\frac{1}{n})$ on typical graphs from $G(n,p)$, is too large to obtain the optimal rate by simply applying the Laplace mechanism. Such a strategy would lead to error $\frac{1}{\varepsilon n}$, instead of $\frac{1}{\varepsilon n^{3/2}}$.
@@ 1.3 Related Work
Privacy in the context of black-box functions was first explicitly considered by [JR13]. They connect the claimed sensitivity setting with the concept of local filter from the sublinear algorithms literature. This line of work [JR13, AJMR14, LLRV25] constructs privacy wrappers from local filters for Lipschitz functions. Their constructions query the function $f$ on inputs obtained by both insertions and deletions. Their query complexity depends on the universe size, and the analyst’s sensitivity bound must allow for the insertion of arbitrary outliers in the dataset.
@@
One can interpret the information-theoretic results of [SCV18] as a generic transformation that takes a function $f$ and produces a version that is robust in the neighborhood of an input $x$ whenever $x$ is resilient with respect to the original function $f$. However, we are not aware of a generic way to transform that into a differentially private mechanism using previous work, without still needing to explore the values of $f(z)$ for all sets $z$ that differ from $x$ in $\lambda$ insertions and deletions. We make the [SCV18] transformation explicit and discuss its consequences in Appendix E.
@@
Dwork and Lei [DL09] and Smith [Smi11] give generic transformations that create differentially private versions of statistical estimators. These results are incomparable to ours, since they rely on specific properties of the estimators, such as asymptotic normality and low bias.
@@ 2 Preliminaries
Definition 2.1 (Neighboring sets, differential privacy [DMNS16]).
@@
When $\delta=0,$ this guarantee is referred to as pure differential privacy; the guarantee with $\delta>0$ is called approximate differential privacy.
@@
Fact 2.1 (Laplace mechanism [DMNS16]).
@@
Fix $\varepsilon>0$ and $\delta\in(0,1)$. Let ${\mathcal{A}}$ be an algorithm and ${\mathcal{M}}$ be an $(\varepsilon,\delta)$-DP mechanism. Then the algorithm that, on input $x$, runs ${\mathcal{A}}$ on the output of ${\mathcal{M}}(x)$, is $(\varepsilon,\delta)$-DP.
@@ 3 Privacy Wrappers with Automated Sensitivity Detection
For every universe $\mathcal{U}$, privacy parameters $\varepsilon>0$ and $\delta\in[0,1)$, error probability $\beta\in[0,1)$, and finite range $\mathcal{Y}\subset\mathbb{R}$ with size $k\stackrel{{\scriptstyle{\mbox{\tiny def}}}}{{=}}|\mathcal{Y}|$, there exist a $\lambda$-down-local privacy wrapper $\mathcal{W}$ such that for every function $f:\mathcal{U}^{*}\to\mathcal{Y}$ and dataset $x\in\mathcal{U}^{*}$, with probability at least $1-\beta$,
@@ 3.1 Shifted Inverse Mechanism: Promised Monotone Functions
In this section, we describe a privacy mechanism that works under a promise that the black-box function $f$ is monotone. It can be viewed as a privacy wrapper under the promise, but we do not call it a privacy wrapper to stress that, in contrast to our privacy wrappers, this mechanism is not private when the promise is broken. Our mechanism is a novel variant [Ste23] of the Shifted Inverse (ShI) mechanism of Fang, Dong, and Yi [FDY22]. The original ShI mechanism satisfies pure differential privacy for all monotone functions $f$. Our variant gets better dependence on the size of the range of function $f$ at the expense of providing only approximate differential privacy.
@@
Now we present our generalization of the ShI mechanism from [FDY22]. We begin by defining the inverse loss function: Given a function $f:\mathcal{U}^{*}\to\mathbb{R}$, we define $\ell^{f}:\mathcal{U}^{*}\times\mathbb{R}\to[0,\infty]$ by [^4] There is an annoying technicality: If $y<f(\emptyset)$, then $\ell^{f}(x,y)=\min\emptyset=+\infty$.
@@
The advantage of the inverse loss function is that it has low sensitivity, even when $f$ has high sensitivity. However, this only holds when $f$ is monotone. The following lemma encapsulates an elegant insight of Fang et al. [FDY22].
@@
Observe that $g(x,j)=1\iff\ell^{f}(x,y_{j})\geq\lambda+1$. That is, the exact value of $\ell^{f}(x,y)$ does not matter past the threshold $\lambda+1$. So we do not need to compute the exact value of $\ell^{f}(x,y)$ in this case. Therefore, we compute $g(x,j)$ using only the values of $f(s)$ on $s\in{\mathcal{N}}^{\downarrow}_{\lambda}(x)$. In symbols, $g(x,j)$ is the maximum of 0 and the following expression:
11 11 $-\frac{1}{\lambda+1}\ell^{f}(x,y_{j})=1-\frac{1}{\lambda+1}\min\{|x\setminus s|:s\subseteq x,f(s)\leq y\}$
$=1-\frac{1}{\lambda+1}\min\left(\{\lambda+1\}\cup\{|x\setminus s|:s\subseteq x,f(s)\leq y\}\right)$
$=1-\frac{1}{\lambda+1}\min\left(\{\lambda+1\}\cup\{|x\setminus s|:s\in{\mathcal{N}}^{\downarrow}_{\lambda}(x),f(s)\leq y\}\right).\qed$
3.2 Sens-o-Matic: A Wrapper for General Functions